20. The average annual premium for automobile insurance in the United States is $1503 Insure.com website, March 6, 2014). The following annual premiums (S) are representa- tive of the websites findings for the state of Michigan. 1905 2725 2677 2600 2962 2184 2332 3112 2545 2525 2370 2545 2529 2442 2312 2981 2627 2857 2675 2115no calculators please! need to know how to work it out the long way, thank you!
Answer #1

a) From the given data,

Premimum (Xi) (Xi-Xbar)^2
1905 417316
2725 30276
2677 15876
2600 2401
2962 168921
2184 134689
2332 47961
3112 314721
2545 36
2525 676
2370 32761
2545 36
2529 484
2442 11881
2312 57121
2981 184900
2627 5776
2857 93636
2675 15376
2115 190096
Total: 51020 1724940

Sample Mean = 51020/20 = 2551

Sample variance (s2) = 1724940 / (20-1) = 90786.31579

SD (s) = 301.3077

a) The point estiamte of the mean annual automobile insurance premium in Michigan is sample mean = $2551

b)

The 95% confidence interval for the mean annual automobile insurance premium in Michigan is
(xbar - tcrit * s/sqrt(n), xbar + tcrit * s/sqrt(n)) = (2551 - 2.093*301.3077/sqrt(20), 2551 + 2.093*301.3077/sqrt(20)) = (2409.983667,2692.016333)

c)


c) Given The average annula premium for automobile insurance in the US is %1503 is not in the above interval. Thus the annula premium for automobile insurance in the Michigan is not include the nationa average for the US.

The 95% chance that the annula premium for automobile insurance in the US is lies between 2409.983667 and 2692.016333